OpenStudy (anonymous):

In a circle, a 90° sector has area 16π ft2. What is the radius of the circle?

OpenStudy (ajprincess):

Area of a sector \[=\pi r^2*\frac{ Angle~of~the~sector }{ 360 }\] \[16=\pi r^2*\frac{ 90 }{ 360}\] Does that help? @oohnaw